In a group of students, ¾ are male. The proportion of male students who like their curry hot is ⅗ and the proportion of female students who like their curry hot is ¼. One student is chosen at random.
- (i)Find the probability that the student chosen is either female, or likes their curry hot, or is both female and likes their curry hot.[4]
- (ii)Showing your working, determine whether the events 'the student chosen is male' and 'the student chosen likes their curry hot' are independent.[2]
Mark scheme
Marking points from the official mark scheme (May/June 2018). The mark or mark type is shown in brackets.
- [B1] (i) Seen, accept unsimplified P(M ∩ H) = 3/4 * 3/5 = 9/20 (0.45)
- [M1] (i) Numerical attempt at P(F) + P(M ∩ H) = 1/4 + 9/20
- [A1] (i) Correct unsimplified expression = 14/20
- [A1] (i) Correct final answer = 7/10 (0.7) OE
- [M1] (ii) Unsimplified, or better, legitimate numerical attempt at P(M) × P(H) and P(M ∩ H) OR Unsimplified, or better, numerical attempt at P(H) and P(M ∩ H), P(M)
- [A1] (ii) Descriptors P(M∩ H) and P(M) × P(H) seen, correct numerical evaluation and comparison, conclusion stated OR Descriptors P(M ∩ H), P(H) and P(M) OR P(M|H) and P(M) seen, numerical evaluation and comparison, conclusion stated
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →